Find the maximum value of
Max F(x, y) = 4 x y ; X>0 y y so Subject to : x² + y = 1 an 16
Answer #1

F = 43 14 95 The laspangeante feenteon 18 Fri) = Pari) + Agai) F= qx1+a( -) o ⇒ 45 +4)=0 + 20 = 0⇒ 4x + AC%)=0 2I+Act)=0 こ う9From += 1 告+答- ) - ok - (3)=8 | x {J4202 P = [虛22) e-G-28) ++ PC (2) |F= 4位 DDR)-29 f moocf =24